Order the folowing from least to greatest: -3/8, -1, 7/5, 1/3, -8/6, 8/6, 99/100.
Answer by Jk22(389)

-3/8, -1, 7/5, 1/3, -8/6, 8/6, 99/100

find the lcm of the denominators :

factors : 8=2^3, 5, 3, 2*3, 2^2*5^2, hence LCM = 2^3*3*5^2 (=600)

the numerator deviennent :

-225, -600, 840, 200, -800, 800, 594, hence the order is :
-8/6, -1, -3/8, 1/3, 99/100, 8/6, 7/5